# Permissive Information-Flow Analysis for Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) are rapidly becoming commodity components of larger software systems. This poses natural security and privacy problems: poisoned data retrieved from one component can change the model’s behavior and compromise the entire system, including coercing the model to spread confidential data to untrusted components. One promising approach is to tackle this problem at the system level via dynamic information flow (aka taint) tracking. Unfortunately, the traditional approach of propagating the most restrictive input label to the output is too conservative for applications where LLMs operate on inputs retrieved from diverse sources. In this paper, we propose a novel, more permissive approach to propagate information flow labels through LLM queries. The key idea behind our approach is to propagate only the labels of the samples that were influential in generating the model output and to eliminate the labels of unnecessary input. We implement and investigate the effectiveness of two variations of this approach, based on (i) prompt-based retrieval augmentation, and (ii) a $k$-nearest-neighbors language model. We compare these with the baseline of an introspection-based influence estimator that directly asks the language model to predict the output label. The results obtained highlight the superiority of our prompt-based label propagator, which improves the label in more than 85% of the cases in an LLM agent setting. These findings underscore the practicality of permissive label propagation for retrieval augmentation.
I Introduction
@@
One possible approach to address this problem is to rely on the LLM itself for mitigation, for example via introspection of the retrieved inputs or guardrails given in the meta-prompt. However, such defenses can be circumvented with more advanced attacks [10, 11, 12, 13], leading to the undesirable cat-and-mouse game that is common in system security.
Figure 1: Illustration of a label propagator (LP) for large language models (LLMs) with tool-calling capabilities. The goal of the LP is to assign the most suitable label to the output of the LLM. In this instance, we consider labels representing trusted and untrusted sources. A naïve LP assigns the most conservative label to the output, which in this example is untrusted. Our proposed influence-based LP takes into account the influence of each retrieved document and determines that the same output can be obtained by solely relying on trusted documents.
@@
Another approach is to tackle this problem at the system level via dynamic information flow (aka taint) tracking [14, 15, 16]. Taint tracking is a standard technique for enforcing integrity and confidentiality properties, and has been used successfully in many applications, including detecting cross-site-scripting vulnerabilities [17] and privacy leaks in mobile applications [18]. In taint tracking, each piece of data is augmented with labels describing its integrity or confidentiality. Taint is usually propagated conservatively: the output of an operation is labeled as the most restrictive (i.e., most confidential or least trusted) label of its inputs. For example, the output of a function that takes two arguments, one trusted and one untrusted, would be labeled as untrusted. A challenge of applying such label propagation mechanisms to LLMs is that the output label would be the upper bound of all inputs (i.e., the context) used for inference. With LLMs having the ability to retrieve documents from different sources, this can quickly become unnecessarily restrictive, a phenomenon known in the literature as label creep [16].
@@
In this paper, we propose a novel approach to propagate more permissive information-flow labels in LLM-based applications which we will refer to as an influence-based label propagator (LP). The key idea of our approach is to propagate only the labels of the samples that were influential in generating the model’s output—and drop the labels of the inputs that are not required. Specifically, for a given context and fixed tolerance $\lambda$, we identify all subcontexts for which the model achieves utility that is at most $\lambda$ below the utility of the full context. Within those subcontexts, we then select the one with the most permissive label. We prove that, under idealizing assumptions, our algorithm identifies the most permissive label(s) possible. An example of our approach is shown in Figure 1.
@@
To ground our work in a relevant LLM system, we implement and evaluate two different realizations of our label propagator: (i) a RAG-based system [19] where the retrieved documents are provided within the prompt to an autoregressive LLM, and (ii) a $k$NN-LM architecture [20], where the output distribution is computed as a mixture of the distribution of the model and the retrieved documents.
@@
We demonstrate the effectiveness of an influence-based label propagator by evaluating it on three datasets: (i) a synthetic dataset containing specific personal details testing the label propagator’s ability to handle a large number of inputs, (ii) a news article dataset assessing whether the LP is able to handle long free-form natural language, and (iii) a dataset consisting of LLM agent conversations as depicted in Figure 1.
@@
II Problem Setting and Goal
@@
This formulation is sufficiently general to represent many real-world applications of LLMs. For example, in retrieval-augmented generation (RAG) [19, 21], the context contains documents retrieved from the knowledge-base. Alternatively, if an LLM uses plugins or function calls to obtain additional information (e.g., web search, email retrieval, calendar query, etc.) [22], the data returned by the plugins is part of the context.
II-A Information-Flow Labels
@@
II-B Goal: Permissive Label Propagation
@@
III Permissive Label Propagation
In this section, we describe our approach and system for the propagation of permissive information-flow labels from the inputs to the LLM output. The core idea is to propagate only the labels of the documents in the context that were necessary for generating the output. A naïve solution to this problem would be to iterate over all subsets of documents in the context and determine whether they can be removed without significantly affecting the model’s output. However, the computational cost of this approach grows exponentially with the number of documents in the context. In this section, we present a more efficient algorithm.
Our key observation is that it is sufficient to identify the labels that improve over the full context’s label, and consider only their corresponding subcontexts. As the label lattices that occur in practice are often small (e.g. secret vs public, trusted vs untrusted), this leads to a solution that is both practical and optimal under certain monotonicity assumptions.
III-A $\lambda$-similar Labels
@@
III-B Computing $\lambda$-similar Labels
@@ III-C Correctness
As utility functions are not necessarily monotonous [26], Algorithm 1 is a heuristic in practice. In Section V, we evaluate how closely it matches the statement in Proposition 1.
III-D A System for Label Propagation
We briefly discuss how to integrate Algorithm 1 into existing model architectures and systems. For this, we consider two architectures for augmenting language models with retrieved information: prompt-based retrieval augmentation [19, 21] and $k$NN language models [20].
@@ III-E Computational Cost
However, the number of LLM queries is always bounded by the size of the lattice. When the lattice forms a totally ordered set (e.g. two-element lattices distinguishing between trusted vs untrusted or confidential vs public data), Algorithm 1 stops as soon as it finds a subcontext whose utility drops below a $\lambda$ difference w.r.t. the utility of the full context. For richer lattices describing more fine-grained security policies, Algorithm 1 visits a small subset of the lattice in typical queries, either because only a few labels are represented in the context or because the utility drops below the tolerance $\lambda$. Therefore, computational costs are not a major concern for several fundamental lattices.
@@
Note that despite an increase in the number of LLM calls for label propagation, the cost incurred per call can be drastically reduced utilizing common inference optimizations implemented in production-ready model serving backends [27, 28]. For instance, compute-bound prompt processing in Transformers can be amortized across calls sharing a common prompt prefix by reusing a KV cache. For a totally ordered lattice, appending retrieved documents at the end of the prompt sorted by their labels will result in no additional prompt processing costs in subsequent LLM calls because their prompts are obtained by peeling off documents at the end of the prompt.
While caching can save computation in processing prompt tokens, it does not impact the efficiency of the memory-bound decoding phase. However, we can use an algorithm-specific optimization to decode only a subset of the tokens that a naïve implementation of Algorithm 1 would decode. Since we use negative perplexity as the utility metric, during the decoding phase of an LLM call, we can keep a running calculation of the cumulative likelihood of the tokens decoded so far and stop decoding as soon as the utility drops below a difference $\lambda$ of the utility of the full context.
IV Evaluation Setup
@@
Figure 4: Sample documents and QA pairs from the synthetic key-value dataset.
@@
Figure 5: Sample documents and QA pairs from the news article dataset.
As this is a novel problem setup, we first define a robust evaluation scheme that we use to evaluate our label propagators and compare them against an introspection baseline. Our goal is to understand the performance of our label propagator in correctly identifying and propagating minimal labels. For this, we construct three datasets that enable us to evaluate different aspects of the system. The evaluation results are presented in Section V.
IV-A Research Questions
@@
Research questions RQ1 and RQ2 focus on the performance of Algorithm 1 directly. Research question RQ3 focuses on the quality of the final output of the end-to-end system introduced in Section III-D.
IV-B Evaluation Metrics
@@
The system for label propagation described in Section III-D regenerates the output based on the inferred label. While the definition of $\lambda$-similarity guarantees that under the subcontext of the inferred label the original output is almost as likely as in the full context, this does not necessarily guarantee that the regenerated output has good utility w.r.t. to other metrics. To quantify potential deviations, we measure the alignment of the different model outputs w.r.t. the target specified in the dataset i.e., $y^{\star}$ using ROUGE-L [29] metric, which is based on the Longest Common Subsequence (LCS) [30] computation and has been commonly used to estimate translation quality in the past.
@@ IV-E Datasets
We use the target response in the dataset as the completion $y$ for Algorithm 1, i.e., we do not rely on the model for generating the output unless mentioned otherwise. This is motivated by the fact that the target label computed in the dataset is only correct w.r.t. the target response. Therefore, an incorrect completion $y$ from the model would render the ground-truth label set $\Lambda^{\star}$ incorrect. We quantify the implications of this decision in Section V-C, where we compare the difference in performance between the dataset target and the model-generated output.
@@
Figure 6: An example interaction from the tool calling LLM case study, where the user query triggers a web search. The items returned by the tool have different labels. The aim of our label propagator is to identify the label for the final model output. Since the high-reliability output is sufficient to answer the query in this case, the label propagator can upgrade the label of the output to high reliability instead of the naive baseline of propagating the low-reliability label.
@@
V Results
V-A Synthetic key-value dataset
@@
TABLE I: Results on the synthetic key-value dataset assuming a perfect recall retriever, which always retrieves the relevant items while also retrieving some irrelevant ones, with person ID as the key, and person SSN and DoB as values. $k$NN-LM prediction uses $\gamma=0.5$. We report mean and one standard deviation for macro-averaged precision and recall.
Figure 7: Exact match accuracy vs. model size on the synthetic key-value dataset. We see no impact of scale on the RAG-based label propagator, a minor impact on the kNN-LM-based label propagator, and a significant impact of scale with introspection.
In this case, we assume a perfect recall retriever (all relevant documents are present), albeit lower precision (the rest of the documents out of the total limit of 14 are filled by adding irrelevant documents). For the introspection baseline, we use one-shot learning based on the first example in the dataset and use the remaining 63 questions for evaluation. Table I summarizes our results.
The prompt-based propagator achieves an exact match accuracy of over 85% and precision and recall of over 90%. This means that, in more than 85% of the cases, the label search is able to identify the correct influential subcontexts out of a set of 16k possibilities without making a single mistake. Furthermore, the RAG-based label search significantly outperforms both the $k$NN-LM-based label search as well as the introspection baseline.
We plot the exact match accuracy of the label search on the key-value dataset w.r.t. the number of parameters in Figure 7. The figure highlights that model size does not strongly correlate with the performance of the RAG-based label propagator, as evident by the near-horizontal line. Consequently, this enables the use of significantly smaller models for effective label propagation on the outputs of larger models.
@@
Sensitivity of hyperparameters. The perplexity tolerance $\lambda$ measures the acceptable loss in model utility when removing a subset of documents. The choice of $\lambda$ allows to trade-off model utility and the restrictiveness of the inferred label.
We illustrate the role of the perplexity tolerance $\lambda$ on the resulting change in the exact match performance for our RAG-based label propagator in Figure 8. The figure indicates that the model performance is very sensitive to the choice of perplexity tolerance i.e., either very high or very low values significantly hamper model performance. Furthermore, the optimal hyperparameters are consistent across model scales. However, larger models, being better at language modeling, outperform their smaller counterparts at lower values of $\lambda$.
Although the influence-based label propagator performs very well on its own, we have found that pruning irrelevant labels can elevate the performance further. This is particularly the case for large lattices with a partial order. We introduce a Shapley-value-based heuristic pruning technique in Appendix -A that can marginally improve the performance from an exact match accuracy of 81% to 86%. The results reported in Table I include this Shapley-value-based heuristic.
V-B News article dataset
@@
TABLE II: Results on the news article dataset with prompt-based label search. The perfect retriever row assumes that the retriever always retrieves the relevant items while also retrieving some irrelevant ones randomly. The realistic retriever uses a cosine similarity-based nearest-neighbour search and may not retrieve all relevant items, thereby reducing the label improvement metric of even a perfect label search.
@@
Table II summarizes the results of the label propagator on the news article dataset. For this dataset, the label propagator achieves slightly worse performance on all metrics in comparison to the synthetic key-value dataset due to the lower quality of the dataset (automated generation of QA pairs by GPT-4 [1]). Furthermore, we see a reduction in the exact match accuracy of about $10\%$ when using a realistic retriever in contrast to a perfect retriever. This is due to the retriever sometimes failing to retrieve the relevant pieces of information, ultimately leading to a mismatch with the ground-truth label.
Due to the presence of total order on the labels, we report label improvement and missed labels for this dataset. Even when a perfect retriever is used, a naïve label propagator would be overly conservative in 647 out of 3465 examples where label improvement is possible. In contrast, our influence-based label propagator is able to assign the correct label in $\sim 57\%$ of these cases. On the other hand, our label propagator suggested a more permissive label but missed a label from the ground-truth in $\sim 23\%$ of the cases. However, as explained in Section III-D, the safety property still holds in these cases because the system regenerates the output using only the articles with the more permissive label (although this means that the regenerated output may differ from the original output, as we quantify in the next section).
@@
Sensitivity of hyperparameters. We highlight the sensitivity to changes in the perplexity threshold $\lambda$ on the news article dataset in Figure 9. Similar to the synthetic key-value dataset, we see similarities across the two model scales, and a higher tolerance of the larger model to lower values of $\lambda$. However, we see a significant rise in optimal perplexity tolerance in contrast to the synthetic dataset due to the higher difficulty of the responses.
V-C LLM agent dataset

### v2

@@ Abstract
Large Language Models (LLMs) are rapidly becoming commodity components of larger software systems. This poses natural security and privacy problems: poisoned data retrieved from one component can change the model’s behavior and compromise the entire system, including coercing the model to spread confidential data to untrusted components. One promising approach is to tackle this problem at the system level via dynamic information flow (aka taint) tracking. Unfortunately, this approach of propagating the most restrictive input label to the output is too conservative for applications where LLMs operate on inputs retrieved from diverse sources.
In this paper, we propose a novel, more permissive approach to propagate information flow labels through LLM queries. The key idea behind our approach is to propagate only the labels of the samples that were influential in generating the model output and to eliminate the labels of unnecessary inputs. We implement and investigate the effectiveness of two variations of this approach, based on (i) prompt-based retrieval augmentation, and (ii) a $k$-nearest-neighbors language model. We compare these with a baseline that uses introspection to predict the output label. Our experimental results in an LLM agent setting show that the permissive label propagator improves over the baseline in more than 85% of the cases, which underscores the practicality of our approach.
1 Introduction
Figure 1: Illustration of a label propagator (LP) for large language models (LLMs) with tool-calling capabilities. The goal of the LP is to assign the most suitable label to the output of the LLM. In this instance, we consider labels representing trusted and untrusted sources. A naïve LP assigns the most conservative label to the output, which in this example is untrusted. The LP we design takes into account the influence of each retrieved document and determines that the same output can be obtained by solely relying on trusted documents.
@@ 1 Introduction
One possible approach to address this problem is to rely on the LLM itself for mitigation, for example via introspection of the retrieved inputs or guardrails given in the meta-prompt. However, such defenses can be circumvented with more advanced attacks [54, 52, 23, 38], leading to the undesirable cat-and-mouse game that is common in system security. A promising alternative is to tackle this problem at the system level via dynamic information-flow (aka taint) tracking. Information-flow tracking is a standard technique for enforcing integrity and confidentiality properties [7, 5, 35], and has been used successfully in many applications, including detecting cross-site-scripting vulnerabilities [43], privacy leaks in mobile applications [12] and recently to LLM-based systems [48, 47, 2].
@@
In information-flow tracking, each piece of data is augmented with labels describing its integrity or confidentiality. Labels are usually propagated conservatively: the output of an operation on data is labeled as the most restrictive (i.e., most confidential or least trusted) label of its inputs. For example, the output of a function that takes two arguments, one trusted and one untrusted, would be labeled as untrusted. A challenge of applying such label propagation mechanisms to LLMs is that the output label would be the upper bound of all inputs (i.e., the context) used for inference. With LLMs having the ability to retrieve documents from different sources, this can quickly become unnecessarily restrictive, a phenomenon known in the literature as label creep [35].
@@
In this paper, we propose a novel approach to propagate more permissive information-flow labels in LLM-based applications which we will refer to as an influence-based label propagator (LP). The key idea of our approach is to propagate only the labels of the samples that were influential in generating the model’s output—and drop the labels of the inputs that are not. Specifically, for a given context and fixed tolerance $\lambda$, we identify all subcontexts for which the model achieves utility that is at most $\lambda$ below the utility of the full context. Within those subcontexts, we then select the one with the most permissive label. We prove that, under idealizing assumptions, our algorithm identifies the most permissive label(s) possible. An example of our approach is shown in Figure 1.
@@
To ground our work in a relevant LLM system, we implement and evaluate two different realizations of our label propagator: (i) a prompt-based system [22] where the retrieved documents are provided within the prompt to an autoregressive LLM, and (ii) a $k$NN-LM architecture [20], where the output distribution is computed as a mixture of the distribution of the model and the retrieved documents.
@@
We demonstrate the effectiveness of our proposed label propagator by evaluating it on three datasets: (i) a synthetic dataset containing personal details for evaluating the label propagator’s ability to handle a large number of inputs, (ii) a news article dataset assessing whether the LP is able to handle long free-form natural language, and (iii) a dataset consisting of LLM agent conversations as depicted in Figure 1 that mimics the applications used with agent frameworks.
@@
2 Problem Setting and Goal
@@
This formulation is sufficiently general to represent many real-world applications of LLMs. For example, in retrieval-augmented generation (RAG) [22, 16], the context contains documents retrieved from the knowledge-base. Alternatively, if an LLM uses plugins to retrieve external data (e.g., web search, email retrieval, calendar query, etc.) [36], the data returned by the plugins is part of the context.
2.1 Information-Flow Labels
@@
Obtaining labels
Example scenarios where labels are readily available are corporate email servers, which label emails from outside as “external”, productivity suites such as Google Workspace [^2] https://support.google.com/a/answer/9292382 or Microsoft SharePoint [^3] https://learn.microsoft.com/en-us/purview/sensitivity-labels-sharepoint-onedrive-files, which implement document classification labels and RBAC, and corporate search engines, which retrieve data from internal and external sources and label them accordingly. Tool-augmentation provides another opportunity to integrate labels by labelling inputs and outputs in tool manifests, a one-off effort that can be factored into the design of each tool.
Threat model
We assume that input labels are given and correct. In the most general case an adversary lives at some point in the information flow lattice and can tamper with documents or tools at or below that level. For the integrity and confidentiality lattices, this means the adversary can poison data labeled as untrusted and read data labeled as public. The adversary’s goal is to cause illicit flows, such as data from low integrity labels (poisoned data) affecting output that has high integrity label, or input data labelled as confidential affecting outputs labeled as public.
2.2 Goal: Permissive Label Propagation
@@
3 Permissive Label Propagation
Our core idea is to propagate only the labels of the documents in the context that were necessary for generating the output. A naïve solution to this problem would be to iterate over all subsets of documents in the context and determine whether they can be removed without significantly affecting the output. However, the computational cost of this approach grows exponentially with the number of documents in the context.
We address this issue with the observation that it is sufficient to identify the labels that improve over the full context’s label, and consider only their corresponding subcontexts. As the label lattices that occur in practice are often small (e.g. secret vs public, trusted vs untrusted), this leads to a solution that is both practical and optimal under a certain monotonicity assumption (mentioned in Eq. 3).
3.1 $\lambda$-similar Labels
@@
3.2 Computing $\lambda$-similar Labels
@@ 3.3 Correctness
As utility functions are not necessarily monotonous [39], Algorithm 1 is a heuristic in practice. In Section 5, we evaluate how closely it matches the statement in Proposition 1. Even finding sub-optimal labels can improve over the baseline label propagator.
3.4 A System for Label Propagation
To integrate Algorithm 1 into existing model architectures and systems, we consider two architectures for augmenting language models with retrieved information: prompt-based retrieval augmentation [22, 16] and $k$NN language models [20].
@@ 3.5 Computational Cost
However, the number of LLM queries is always bounded by the size of the lattice. When the lattice forms a totally ordered set (e.g. two-element lattices distinguishing between trusted vs untrusted or confidential vs public data), Algorithm 1 stops as soon as it finds a subcontext whose utility drops below a $\lambda$ difference w.r.t. the utility of the full context. For richer lattices describing more fine-grained security policies, Algorithm 1 visits a small subset of the lattice in typical queries, either because only a few labels are represented in the context or because the utility drops below the tolerance $\lambda$. Therefore, computational costs are not a major concern for several fundamental lattices. We discuss several optimizations in Section 6.3 that reduce the cost of running the system.
@@
4 Evaluation Setup
@@
Figure 4: Sample documents and QA pairs from the synthetic key-value dataset. The question refers to the social security numbers and dates of birth of person 1 and 2. This information can be obtained by accessing documents $A$, $B$, and $C$ or $A$, and $D$, hence the resulting label of $\{ABC,AD\}$.
@@
Figure 5: Sample documents and QA pairs from the news article dataset. The first question can only be answered with access to the LoInt document whereas the second question can be also answered with the HiInt document.
As this is a novel problem setup, we first define a robust evaluation scheme that we use to evaluate our approach and compare against an introspection baseline. Our goal is to understand the performance of our label propagator in correctly identifying and propagating minimal labels.
4.1 Research Questions
@@
Research questions RQ1 and RQ2 focus on the performance of Algorithm 1 directly. Research question RQ3 focuses on the quality of the final output of the end-to-end system introduced in Section 3.4.
4.2 Evaluation Metrics
@@
The system for label propagation described in Section 3.4 regenerates the output based on the inferred label. While the definition of $\lambda$-similarity guarantees that under the subcontext of the inferred label the original output is almost as likely as in the full context, this does not necessarily guarantee that the regenerated output has good utility w.r.t. to other metrics. To quantify potential deviations, we measure the alignment of the different model outputs w.r.t. the target specified in the dataset i.e., $y^{\star}$ using ROUGE-L [24] metric, which is based on the Longest Common Subsequence (LCS) [17] computation and has been commonly used to estimate translation quality in the past.
@@ 4.5 Datasets
We use the target response in the dataset as the completion $y$ for Algorithm 1, i.e., we do not rely on the model for generating the output unless mentioned otherwise. This is motivated by the fact that the target label computed in the dataset is only correct w.r.t. the target response. Therefore, an incorrect completion $y$ from the model would render the ground-truth label set $\Lambda^{\star}$ incorrect. We quantify the implications of this decision in Section 5.3, where we compare the difference in performance between the dataset target and the model-generated output.
@@
Figure 6: An example interaction from the tool calling LLM case study, where the user query triggers a web search. The items returned by the tool have different labels. The aim of our label propagator is to identify the label for the final model output. Since the high-integrity output is sufficient to answer the query in this case, the label propagator can upgrade the label of the output to high integrity instead of the naive baseline of propagating the low-integrity label.
@@
5 Results
5.1 Synthetic key-value dataset
@@
Table 1: Results on the synthetic key-value dataset assuming a perfect recall retriever, which always retrieves the relevant items while also retrieving some irrelevant ones, with person ID as the key, and person SSN and DoB as values. $k$NN-LM prediction uses $\gamma=0.5$. We report mean and one standard deviation for macro-averaged precision and recall.
In this case, we assume a perfect recall retriever (all relevant documents are present), albeit lower precision (the rest of the documents out of the total limit of 14 are filled by adding irrelevant documents). For the introspection baseline, we use one-shot in-context learning [46] based on the first example in the dataset and use the remaining 63 questions for evaluation. Table 1 summarizes our results.
The prompt-based propagator achieves an exact match accuracy of over 85% and precision and recall of over 90%. That is, in more than 85% of the cases, the label search can identify the correct influential subcontexts out of a set of 16k possibilities without error. Furthermore, the prompt based label search significantly outperforms both the $k$NN-LM-based label search as well as the introspection baseline.
Table 1 also compares the exact match accuracy for two different model sizes 7B and 70B parameters. The model size does not strongly correlate with the performance of the prompt based label propagator. Consequently, this enables the use of significantly smaller models for effective label propagation on the outputs of larger models.
@@
Sensitivity of hyperparameters The perplexity tolerance $\lambda$ measures the acceptable loss in model utility when removing a subset of documents. The choice of $\lambda$ allows to trade-off model utility and the restrictiveness of the inferred label.
We illustrate the role of the perplexity tolerance $\lambda$ on the resulting change in the exact match performance for our prompt based label propagator in Figure 7. The figure indicates that the model performance is very sensitive to the choice of perplexity tolerance i.e., either very high or very low values significantly hamper model performance. Furthermore, the optimal hyperparameters are consistent across model scales. However, larger models, being better at language modeling, outperform their smaller counterparts at lower values of $\lambda$.
Although the influence-based label propagator performs very well on its own, we have found that pruning irrelevant labels can elevate the performance further. This is particularly the case for large lattices with a partial order. We introduce a Shapley-value-based heuristic pruning technique in Appendix A that can marginally improve the performance from an exact match accuracy of 81% to 86%. The results reported in Table 1 include this Shapley-value-based heuristic.
5.2 News article dataset
@@
Table 2: Results on the news article dataset with prompt-based label search. The perfect retriever row assumes that the retriever always retrieves the relevant items while also retrieving some irrelevant ones randomly. The realistic retriever uses a cosine similarity-based nearest-neighbour search and may not retrieve all relevant items, thereby reducing the label improvement metric of even a perfect label search.
@@
Table 2 summarizes the results of the label propagator on the news article dataset. For this dataset, the label propagator achieves slightly worse performance on all metrics in comparison to the synthetic key-value dataset due to the lower quality of the dataset (automated generation of QA pairs by GPT-4 [32]). Furthermore, we see a reduction in the exact match accuracy of about $10\%$ when using a realistic retriever in contrast to a perfect retriever. This is due to the retriever sometimes failing to retrieve the relevant pieces of information, ultimately leading to a mismatch with the ground-truth label.
Due to the presence of total order on the labels, we report label improvement and missed labels for this dataset. Even when a perfect retriever is used, a naïve label propagator would be overly conservative in 647 out of 3465 examples where label improvement is possible. In contrast, our influence-based label propagator is able to assign the correct label in $\sim 57\%$ of these cases. On the other hand, our label propagator suggested a more permissive label but missed a label from the ground-truth in $\sim 23\%$ of the cases. However, as explained in Section 3.4, the safety property still holds in these cases because the system regenerates the output using only the articles with the more permissive label (although this means that the regenerated output may differ from the original output, as we quantify in the next section).
@@
Sensitivity of hyperparameters We highlight the sensitivity to changes in the perplexity threshold $\lambda$ on the news article dataset in Figure 8. Similar to the synthetic key-value dataset, we see similarities across the two model scales, and a higher tolerance of the larger model to lower values of $\lambda$. However, we see a significant rise in optimal perplexity tolerance in contrast to the synthetic dataset due to the higher difficulty of the responses.
5.3 LLM agent dataset